OpenStudy (calculusxy):

Choose positive integers m and n, with m < n. Let x = 2mn, y = n^2 −m^2, and z = m^2 + n^2. It so happens that these three positive integers x, y, and z have a special property. What is the property? Can you prove a general result?

OpenStudy (calculusxy):

@agent0smith

OpenStudy (calculusxy):

@skullpatrol @Nnesha @Directrix

OpenStudy (calculusxy):

I had n = 4 and m = 2. So I got: x = 16 y = 12 z = 20

OpenStudy (calculusxy):

I don't know what's so special about these three numbers though except they are the multiplies of 2 and 4?

Directrix (directrix):

Do you know what is special about this triple of numbers: ( 3, 4, 5 ) ?

OpenStudy (calculusxy):

They are Pythagorean Triples! That makes so much more sense!

OpenStudy (calculusxy):

16/4 = 4 12/4 = 3 20/4 = 5

OpenStudy (calculusxy):

What do they mean by proving the "general result" ?

Directrix (directrix):

Correct. Now, your task is to prove a general result for that with the variables that were given.

OpenStudy (calculusxy):

I still don't understand.

OpenStudy (agent0smith):

x = 2mn, y = n^2 −m^2, and z = m^2 + n^2 Can't you just show that \(\large x^2 + y^2 = z^2 \) ?

OpenStudy (agent0smith):

This looks like it'd be true to me, from some mental checking\[\large (2mn)^2 + (n^2 - m^2)^2 = (m^2 + n^2)^2\]